OpenStudy (gavin39):

Primo car rental agency charges ​$32 per day plus ​$0.35 per mile. Ultimo car rental agency charges ​$34 per day plus ​$1.00 per mile. Find the daily mileage for which the Ultimo charge is twice the Primo charge.

OpenStudy (gavin39):

I give medals!!

OpenStudy (gavin39):

@563blackghost

OpenStudy (ajspeller):

32 + 0.35x = 2(34 + x) Solve for x

OpenStudy (gavin39):

Not to aware how to go about that

OpenStudy (ajspeller):

I made a mistake in my earlier post. Follow this post please. Primo is represented by the equation y = $32 + 0.35x Ultimo is represented by the equation y = $34 + 1x So, we want to find out when Ultimo is twice Primo Ultimo = twice(Primo) Ultimo = 2(Primo) $34 + 1x = 2($32+ 0.35x) Then you solve for x.